OpenStudy (anonymous):

Calculate the value of x that makes the expression true; round the answer to the nearest ten-thousandth. ex – 5 = 8 Select one: a. 9.2564 b. 6.5942 c. 5.9642 d. 2.5649

OpenStudy (anonymous):

@TylerD

OpenStudy (kc_kennylau):

Add 5 to both sides then log both sides

OpenStudy (tylerd):

ex=13

OpenStudy (tylerd):

x=13/e

OpenStudy (tylerd):

except that doesnt actually work

OpenStudy (kc_kennylau):

I think he means \(e^x\) instead of \(e\times x\)

OpenStudy (tylerd):

oh

OpenStudy (mokeira):

@kc_kennylau i think so too

OpenStudy (tylerd):

\[\ln e^x = \ln13\] \[x=\ln13\]

OpenStudy (tylerd):

lne = 1

OpenStudy (tylerd):

and when you have an exponent you move it so